OpenStudy (kalebj):

WILL MEDAL AND FAN What is 4.86 × 104 written in standard form? A. 0.0000486 B. 0.000486 C. 48,600 D. 486,000

OpenStudy (kalebj):

@sweetburger

OpenStudy (kalebj):

@mathmate

OpenStudy (kalebj):

@Ms-Brains

MsBrains (ms-brains):

Any ideas? :)

OpenStudy (kalebj):

Thinking b ☺☻

OpenStudy (kalebj):

@DallasCowboys88

OpenStudy (mathmate):

4.86 × 104 = 505.44 Your question is probably 4.86 × 10^4=\(4.86 × 10^4 \) If the exponent of 10 is positive, move the decimal that number of places to the \(right\). Otherwise move to the left. Example: \(2.5\times 10^2=2.5\times 100=250\)

OpenStudy (kalebj):

Sooo its B.

OpenStudy (kalebj):

No no wait its A.